Calculate the voltage of a battery connected to a parallel plate area of 2.00×10-2 m 2 and a plate separation of 2.0×10-³ m if the charge stored is 4.00×10-¹² C



1
Expert's answer
2022-05-18T08:31:15-0400
"V=q/C=qd/(\epsilon_0 A)\\
=\frac{4.00*10^{-12}*2.0*10^{-3}}{8.85*10^{-12}*2.00*10^{-2}}=4.52*10^{-2}\rm\: V"